Shift summary for Aug. 7th (Owl)

Elliott Cheu (elliott@physics.arizona.edu)
Sat, 07 Aug 1999 05:05:16 -0700

Shift Summary for OWL shift, 07/AUG/99

On Shift: ECC, MB

Hours of beam available at all intensities: 6.5/7.0
Hours of beam available at nominal intensity: 6.5/7.0
Hours of E832 data (special runs not included): 6.5/7.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: kfrends

========================================================================
Runs ended this shift: None

Tapes brought to FCC this shift: None
(including tapes for online split)

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:

E832 datataking, Beam is intermittent with lots of aborts.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14133 00:30 Notice that L1 source latch distributions
look screwed up. Probably related to AZLAT
removal.

1:05 BOA above 1.5. Called MCR and it went down.
2:45 BOA above 1.4. Called MCR again and it goes down.

3:30 ktevana errors. kfrend, reset kumquats,
kfrend.

3:40 Notice that L3 process on ktev3 had died at 2:25.
Doesn't seem to affect livetime. This process
died in T3EOP.

4:20 Laser crate full. kfrend. reset pipeline. kfrend.

5:15 ktevana errors. kfrend.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

L3_driver_mon_832 remade to fix HCC inefficiency alarms. This will
take effect on next run.

Please brute force L3 processes. This will use the new monitoring code and
will fix the dead L3 process on ktev3.

If beam disappears for more than an hour, please run a laser scan.